Find four consecutive even integers whose sum is 340: Find four consecutive even integers whose sum is 340
Answer by Alan3354(1431)

Find four consecutive even integers whose sum is 340
-----
Since it spece even, use 2n.
2n+(2n+2)+(2n+4)+(2n+6) = 340
8n + 12 = 340
8n = 328
2n = 82 (The 1st number)
So it's 82+84+86+88 = 340
-------------------------
Using N would work the same, as there are not 4 odd integers that do that.